# Hermite trace polynomials and chaos decompositions for the Hermitian Brownian motion

### v2

@@ 4. Fock space
the vector space quotient under this action, which may be identified with the fixed point subspace, the image of the direct sum of the projections
@@
On $\mathbb{C}{(0)}\oplus\bigoplus_{n=1}^{\infty}\left(\mathbb{C}[S_{0}(n)]\otimes\mathcal{H}^{\otimes n}\right)$, we have the canonical inner product
@@
where $\chi_{0}$ is the canonical trace on $\mathbb{C}[S_{0}(n)]$. On this space, define the operator
(5) $\mathcal{K}_{q}\left(\alpha\otimes F\right)=n!\sum_{\beta\in S_{0}(n)}\chi_{q}[\alpha\beta^{-1}]\beta\otimes F.$
Note that $\mathcal{K}_{0}$ is the identity operator. It is easy to check that for $q\in\mathcal{Z}$, $\mathcal{K}_{q}$ is a positive semi-definite operator. Moreover,
@@
so $\mathcal{K}_{q}$ restricts to the subspace $\overline{\mathcal{TP}}(\mathcal{H})$. We thus define the inner product on $\mathbb{C}{(0)}\oplus\bigoplus_{n=1}^{\infty}\left(\mathbb{C}[S_{0}(n)]\otimes\mathcal{H}^{\otimes n}\right)$ by
(6) $\begin{split}\left\langle\alpha\otimes F,\beta\otimes G\right\rangle_{q}&=\left\langle\alpha\otimes F,\mathcal{K}_{q}P(\beta\otimes G)\right\rangle_{f}\\ &=\delta_{n=k}\sum_{\sigma\in S(n)}\chi_{q}(\alpha\sigma\beta^{-1}\sigma^{-1})\left\langle F,U_{\sigma}G\right\rangle_{\mathcal{H}^{\otimes n}}.\end{split}$
For this inner product, $P_{n}$ is an isometric projection. The resulting inner product on $\overline{\mathcal{TP}}(\mathcal{H})$ takes a simpler form
@@
and so the 00-inner product on $\overline{\mathcal{TP}}(\mathcal{H})$ is positive definite.
@@
In [GM02], Guţă and Maassen considered a general Fock space construction. Let $\mathcal{H}$ and $\left\{V_{n}:n\in\mathbb{N}\right\}$ be Hilbert spaces, such that $S(n)$ acts unitarily on $V_{n}$. Then one can define $V_{n}\otimes_{s}\mathcal{H}^{\otimes n}$ as the fixed point subspace of the action
@@
where $\mathcal{F}_{f}(\mathcal{H})$ is the full Fock space of $\mathcal{H}$.
@@
Note that $\mathcal{TP}_{q}(\mathbb{C})$ is not $\mathcal{TP}_{q}$ from Proposition 3.5, but its symmetrized version,
$\mathcal{TP}_{q}(\mathbb{C})=\bigoplus_{n=0}^{\infty}Z(\mathbb{C}[S_{0}(n)]:\mathbb{C}[S(n)])/\mathcal{N}_{gr,q}.$
@@
    If $\mathcal{H}$ is infinite-dimensional, the linear span of the elements of the form $(\alpha\otimes(h_{1}\otimes\ldots\otimes h_{n}))$ for mutually orthogonal $h_{1},\ldots,h_{n}$ is dense in $\mathcal{F}_{q}(\mathcal{H})$.
@@
Part (b) follows from the fact that for infinite-dimensional $\mathcal{H}$, the linear span of the elements of the form $h_{1}\otimes\ldots\otimes h_{n}$ for mutually orthogonal $h_{1},\ldots,h_{n}$ is dense in $\mathcal{H}^{\otimes n}$. ∎
@@
$\Delta(\Xi^{n})=\left\{\mathbf{u}\in\Xi^{n}:u(1)\leq u(2)\leq\ldots\leq u(n)\right\},$
[…]
$\begin{split}Z(\mathbb{C}[S_{0}(n)]:\pi)&=Z(\mathbb{C}[S_{0}(n)]:S(I_{1})\times\ldots\times S(I_{k})).\end{split}$
@@
    (9) $\overline{\mathcal{TP}}(\mathcal{H})=\bigoplus_{n=0}^{\infty}\bigoplus_{\mathbf{u}\in\Delta(\Xi^{n})}Z(\mathbb{C}[S_{0}(n)]:\ker(\mathbf{u}))\otimes_{s}(\xi_{u(1)}\otimes\ldots\otimes\xi_{u(n)})$
    which is orthogonal with respect to any $q$-inner product.
@@
    where $\eta_{\mathbf{u}}\in Z(\mathbb{C}[S_{0}(n)]:\ker(\mathbf{u}))\cap\mathbb{C}[S_{0}(n)]_{\leq N}$, and
    $\left\langle A,A\right\rangle_{q}=\sum_{n=0}^{\infty}\sum_{\mathbf{u}\in\Delta(\Xi^{n})}\ker(\mathbf{u})!\chi_{q}[\eta_{\mathbf{u}}^{\ast}\eta_{\mathbf{u}}]<\infty,$
@@
    For each $n$, for sufficiently large $N$, $\mathbb{C}[S_{0}(n)]_{\leq N}\otimes_{s}\mathcal{H}^{\otimes n}$ is complete with respect to the $\frac{1}{N}$-norm.
@@
For orthogonality, we observe that if $\mathbf{u},\mathbf{v}\in\Delta(\Xi^{n})$,
@@
Part (b) follows from the fact that each subspace $Z(\mathbb{C}[S_{0}(n)]:\ker(\mathbf{u}))\otimes(\xi_{u(1)}\otimes\ldots\otimes\xi_{u(n)})$ is finite dimensional, and thus closed.
@@
Let $\mathcal{H}=\mathbb{C}$, so that $\mathcal{TP}(\mathbb{C})=Z(\mathbb{C}[S_{0}(n)]:\mathbb{C}[S(n)])$. Then $\left\{\chi^{\lambda^{\prime}:\lambda}:\lambda\in\operatorname{\mathrm{Par}}(n),\lambda^{\prime}=\lambda+\square\right\}$ are orthogonal and span $\mathcal{TP}(\mathbb{C})$. Moreover,
@@
Every element in the $n$’th component of $\overline{\mathcal{TP}}(\mathcal{H}_{\mathbb{R}})$ is equivalent to a unique element of the form
@@ 5. The operator algebra
We now define a star-algebra structure on $\mathcal{TP}(\mathcal{H}_{\mathbb{R}})$, and eventually on $\overline{\mathcal{TP}}(\mathcal{H}_{\mathbb{R}})$. To distinguish the elements of the algebra from the corresponding elements of the inner product space, we will denote $\alpha\otimes_{s}F$ by $\mathnormal{W}\left(\alpha\otimes_{s}F\right)$ to emphasize its interpretation as a Wick product. Note that this identification differs from the one in Section 3.2.
@@
More generally, for $\pi\in S(n)$ with the cycle structure $\pi\in\operatorname{\mathcal{P}}_{1,2}(n)$,
@@
The (tensor) contraction is not a bounded operator and so does not extend to the Hilbert space tensor product $\mathcal{H}^{\otimes n}$. But if $\pi\in\operatorname{\mathcal{P}}(n,k)$, it is easy to check that
$C_{\pi}:\mathcal{H}^{\otimes n}\times\mathcal{H}^{\otimes k}\rightarrow\mathcal{H}^{\otimes(n+k-2\ell)}$
is a contraction.
@@
$\tilde{\sigma}=P^{[0,n]\setminus\mathrm{supp}\left(\pi\right)}_{[0,n-2\ell]}\sigma P_{[0,n]\setminus\mathrm{supp}\left(\tilde{\pi}\right)}^{[0,n-2\ell]}\in S(n-2\ell).$
@@
Then $\mathrm{T}$ is also well-defined on $\mathcal{TP}(\mathcal{H}_{\mathbb{R}})$. Note that we cannot in general extend it to $\overline{\mathcal{TP}}(\mathcal{H}_{\mathbb{R}})$.
@@
For $\alpha\in S_{0}(n)$,
$\begin{split}\mathnormal{W}\left(\alpha\otimes F\right)=\mathnormal{T}\left(e^{-\mathcal{L}}(\alpha\otimes F)\right)&=\sum_{k=0}^{\infty}(-1)^{k}\frac{1}{k!}T\left(\mathcal{L}^{k}(\alpha\otimes F)\right)\\ &=\sum_{\pi\in\operatorname{\mathcal{P}}_{1,2}(n)}(-1)^{n-\left|\pi\right|}\mathnormal{T}\left(C_{\pi}(\alpha\otimes F)\right).\end{split}$
@@
The multiplication on $\mathcal{TP}(\mathcal{H})$ is well defined, and
$(\mathnormal{T}\left(\alpha\otimes_{s}F\right)\mathnormal{T}\left(\beta\otimes_{s}G\right))^{\ast}=\mathnormal{T}\left(\beta\otimes_{s}F\right)^{\ast}\mathnormal{T}\left(\alpha\otimes_{s}G\right)^{\ast}.$
@@
For $\sigma,\tau\in S(n)$, taking $\rho=\sigma\sigma_{n,k}^{-1}\tau\sigma_{n,k}$
@@
For $q\in\mathcal{Z}\setminus\left\{0\right\}$, $\mathcal{N}_{vs,q}$ is an ideal for multiplication in Definition 5.7, and is preserved by each $C_{\pi}$. Consequently, $\mathcal{L}$ is defined as a map on $\mathcal{TP}_{q}(\mathcal{H}_{\mathbb{R}})$, and $\mathcal{L}_{n,k}^{(\ell)}$ on the appropriate quotient. Also, $\mathnormal{T}\left(\eta\right)$ is well-defined for $\eta\in\mathcal{TP}_{q}(\mathcal{H}_{\mathbb{R}})$.
@@
We obtain an extension of the familiar factorization property of Hermite polynomials.
@@
    for $\eta\in\mathbb{C}[S_{0}(n)]$. Then for such $\eta$,
@@
In particular, anything of lower degree is in the sum of eigenspaces with eigenvalues $10,1,\ldots,n-$ 1. It follows that specifying the leading term of an eigenfunction with a given eigenvalue determines it. ∎
Notation 5.13.
If $C_{\pi}(\eta\otimes F)$ is a scalar multiple of ${(0)}$ (which is the identity for the algebra), we will identify it with a scalar.
@@
Define a unital linear functional on $\bigoplus_{n=0}^{\infty}\mathbb{C}[S_{0}(n)]\otimes\mathcal{H}_{\mathbb{R}}^{\odot n}$ by
$\varphi\left[\mathnormal{W}\left(\alpha\otimes F\right)\right]=0,\quad\varphi\left[\mathnormal{W}\left({(0)}\right)\right]=1.$
Then
@@
    In particular, $\ast$ is the adjoint with respect to this inner product, and $\varphi$ is well-defined on the quotient $\mathcal{TP}_{q}(\mathcal{H}_{\mathbb{R}})$.
@@
    $\left\{A:\varphi\left[A^{\ast}A\right]=0\right\}=\mathrm{Span}\left(\mathnormal{W}\left(\zeta\right):\zeta\in\mathcal{N}_{vs,q}\right),$
    and $\varphi$ is faithful on $\mathcal{TP}_{q}(\mathcal{H}_{\mathbb{R}})$.
@@
Moreover each cycle of
@@
intersects $[0,n]$. A computation shows that its restriction to $[0,n]$ is $\sigma^{-1}\alpha\sigma\beta^{-1}$. Therefore
$q^{n-\operatorname{\mathrm{cyc}}_{0}(\sigma^{-1}\sigma_{n,n}(\beta^{-1}\cup\alpha))}=q^{n-\operatorname{\mathrm{cyc}}_{0}(\sigma^{-1}\alpha\sigma\beta^{-1})}=\chi_{q}(\sigma^{-1}\alpha\sigma\beta^{-1})$
@@
Consequently, the star-algebra structure extends to $\overline{\mathcal{TP}}(\mathcal{H})$ and $\overline{\mathcal{TP}}_{q}(\mathcal{H})$.
@@
$\left|\operatorname{\mathcal{P}}_{1,2}(n,k)\right|$ is sequence A086885 in OEIS, and has an easy estimate
$\left|\operatorname{\mathcal{P}}_{1,2}(n,k)\right|=\sum_{\ell=0}^{\min(n,k)}\ell!\binom{n}{\ell}\binom{k}{\ell}=\sum_{\ell=0}^{\min(n,k)}\frac{n!}{(n-\ell)!}\binom{k}{\ell}\leq n^{k}2^{k}.\qed$
@@
For $\alpha_{i}\in S_{0}(n_{i})$, and $F_{i}\in\overline{\mathcal{TP}}(\mathcal{H})$,
@@
Here $\operatorname{\mathcal{P}}_{2}(n_{1},\ldots,n_{k})$ are the inhomogeneous pair partitions [dSCV85].

### v3

@@ 4. Fock space
the direct sum of vector space quotients under these actions, which may be identified with (the direct sum of) the fixed point subspace(s), the image of the direct sum of the projections
@@
On $\mathbb{C}{(0)}\oplus\bigoplus_{n=1}^{\infty}\left(\mathbb{C}[S_{0}(n)]\otimes\mathcal{H}^{\otimes n}\right)$, we have the canonical (“free”) inner product
@@
where $\chi_{0}$ is the canonical trace on each $\mathbb{C}[S_{0}(n)]$. On this space, define the operator
(7) $\mathcal{K}_{q}\left(\alpha\otimes F\right)=n!\sum_{\beta\in S_{0}(n)}\chi_{q}[\alpha\beta^{-1}]\beta\otimes F,$
for $\alpha\in S_{0}(n)$ and $F\in\mathcal{H}^{\otimes n}$. Note that $\mathcal{K}_{0}$ is the identity operator. It is easy to check that for $q\in\mathcal{Z}$, $\mathcal{K}_{q}$ is a positive operator. Moreover,
@@
so $\mathcal{K}_{q}$ restricts to the subspace $\overline{\mathcal{TP}}(\mathcal{H})$. We thus define a (degenerate) inner product on $\mathbb{C}{(0)}\oplus\bigoplus_{n=1}^{\infty}\left(\mathbb{C}[S_{0}(n)]\otimes\mathcal{H}^{\otimes n}\right)$ by
(8) $\begin{split}\left\langle\alpha\otimes F,\beta\otimes G\right\rangle_{q}&=\left\langle\alpha\otimes F,\mathcal{K}_{q}P(\beta\otimes G)\right\rangle_{f}\\ &=\delta_{n=k}\sum_{\sigma\in S(n)}\chi_{q}(\alpha\sigma\beta^{-1}\sigma^{-1})\left\langle F,U_{\sigma}G\right\rangle_{\mathcal{H}^{\otimes n}}\end{split}$
for $F\in\mathcal{H}^{\otimes n}$ and $G\in\mathcal{H}^{\otimes k}$. The map $P_{n}$ is an orthogonal projection for this inner product, but (because of the degeneracy) it is also an isometry. The resulting inner product on $\overline{\mathcal{TP}}(\mathcal{H})$ takes a simpler form
@@
and so the 0-inner product on $\overline{\mathcal{TP}}(\mathcal{H})$ is positive definite.
@@
In [18], Guţă and Maassen considered a general Fock space construction. Let $\mathcal{H}$ and $\left\{V_{n}:n\in\mathbb{N}\right\}$ be Hilbert spaces, such that $S(n)$ acts unitarily on $V_{n}$. Then one can define $V_{n}\otimes_{s}\mathcal{H}^{\otimes n}$ as the fixed point subspace of the action
@@
where $\mathcal{F}_{f}(\mathcal{H})$ is the full Fock space of $\mathcal{H}$. Here $\mathcal{N}_{gr,q}$ is the kernel of the inner product induced by $\chi_{q}$ on the (direct sum of) group algebra(s) $\bigoplus_{n=0}^{\infty}\mathbb{C}[S_{0}(n)]$, while $\mathcal{N}_{vs,q}$ is the kernel of the $q$-inner product on the vector space $\overline{\mathcal{TP}}(\mathcal{H})$.
@@
Note that $\overline{\mathcal{TP}}_{q}(\mathbb{C})$ is not $\mathcal{TP}_{q}$ from Proposition 3.5, but its symmetrized version,
$\overline{\mathcal{TP}}_{q}(\mathbb{C})=\bigoplus_{n=0}^{\infty}Z(\mathbb{C}[S_{0}(n)]:\mathbb{C}[S(n)])/\mathcal{N}_{gr,q}.$
@@
    If $\mathcal{H}$ is infinite-dimensional, the linear span of the elements of the form $\alpha\otimes(h_{1}\otimes\ldots\otimes h_{n})$ for mutually orthogonal $h_{1},\ldots,h_{n}$ is dense in $\mathcal{F}_{q}(\mathcal{H})$.
@@
Part (b) follows from the fact that for infinite-dimensional $\mathcal{H}$, the linear span of the elements of the form $h_{1}\otimes\ldots\otimes h_{n}$ for mutually orthogonal $h_{1},\ldots,h_{n}$ is dense in $\mathcal{H}^{\otimes n}$ (for $\mathcal{H}=L^{2}([0,1],dx)$, this is a well-known fact in stochastic analysis). ∎
@@
Notation 4.7.
Let $\pi=(I_{1},\ldots,I_{k})\in\operatorname{\mathit{Int}}(n)$ be an interval partition of $n$ elements. Denote $S(I_{j})$ the permutations of the set $I_{j}$. We can then identify $S(I_{1})\times\ldots\times S(I_{k})$ with a subgroup of $S(n)$, and so also of $S_{0}(n)$. Its centralizer, which we denote
$\begin{split}Z(\mathbb{C}[S_{0}(n)]:\pi)&=Z(\mathbb{C}[S_{0}(n)]:S(I_{1})\times\ldots\times S(I_{k})),\end{split}$
consists of all elements of $S_{0}(n)$ commuting with this subgroup.
@@
$\Delta(\Xi^{n})=\left\{\mathbf{u}\in\Xi^{n}:u(1)\leq u(2)\leq\ldots\leq u(n)\right\}.$
For $\mathbf{u}\in\Xi^{n}$, denote $\ker(\mathbf{u})$ the interval partition $\pi=(I_{1},\ldots,I_{k})\in\operatorname{\mathit{Int}}(n)$ such that $u(i)=u(j)\Leftrightarrow i\stackrel{{\scriptstyle\pi}}{{\sim}}j$, and $Z(\mathbb{C}[S_{0}(n)]:\pi)$ the centralizer of the corresponding subgroup.
@@
    (11) $\overline{\mathcal{TP}}(\mathcal{H})=\bigoplus_{n=0}^{\infty}\bigoplus_{\mathbf{u}\in\Delta(\Xi^{n})}Z(\mathbb{C}[S_{0}(n)]:\ker(\mathbf{u}))\otimes_{s}(\xi_{u(1)}\otimes\ldots\otimes\xi_{u(n)}).$
    If $q\in\mathcal{Z}$, this decomposition is orthogonal with respect to any $q$-inner product.
@@
    where $\eta_{\mathbf{u}}\in Z(\mathbb{C}[S_{0}(n)]:\ker(\mathbf{u}))$ for $n<N$ and $\eta_{\mathbf{u}}\in Z(\mathbb{C}[S_{0}(n)]:\ker(\mathbf{u}))\cap\mathbb{C}[S_{0}(n)]_{\leq N}$ for $n\geq N$, and
    (12) $\left\langle A,A\right\rangle_{q}=\sum_{n=0}^{\infty}\sum_{\mathbf{u}\in\Delta(\Xi^{n})}\ker(\mathbf{u})!\chi_{q}[\eta_{\mathbf{u}}^{\ast}\eta_{\mathbf{u}}]<\infty,$
@@
    For each $n$, for sufficiently large $N$, $\mathbb{C}[S_{0}(n)]\otimes_{s}\mathcal{H}^{\otimes n}$ is complete with respect to the $\frac{1}{N}$-norm.
@@
For orthogonality, we observe that if $\mathbf{u},\mathbf{v}\in\Delta(\Xi^{n})$ and $\eta,\zeta\in Z(\mathbb{C}[S_{0}(n)]:\ker(\mathbf{u}))$,
@@
Here the second equality holds because for any $\sigma$ exchanging elements of different $I_{i}$ and $I_{j}$, the inner product $\left\langle\xi_{\mathbf{u}},U_{\sigma}\xi_{\mathbf{v}}\right\rangle$ is zero. The third equality follows from the choice that $\zeta$ is in the centralizer.
Part (b) for $\overline{\mathcal{TP}}_{1/N}(\mathcal{H})$ follows from part (a) and Definition 4.4. To extend the decomposition to $\mathcal{F}_{1/N}(\mathcal{H})$, we note that each subspace $Z(\mathbb{C}[S_{0}(n)]:\ker(\mathbf{u}))\otimes(\xi_{u(1)}\otimes\ldots\otimes\xi_{u(n)})$ is finite dimensional, and thus closed, and the norm-squared (12) is the sum of norms-squared on these subspaces.
@@
Let $\mathcal{H}=\mathbb{C}$, so that $\mathcal{TP}(\mathbb{C})=Z(\mathbb{C}[S_{0}(n)]:\mathbb{C}[S(n)])$. Then
$\left\{\chi^{\lambda^{\prime}:\lambda}:\lambda\in\operatorname{\mathrm{Par}}(n),\lambda^{\prime}=\lambda+\square\right\}$
are orthogonal and span $\mathcal{TP}(\mathbb{C})$. Moreover,
@@
Every element in the $n$’th component of $\overline{\mathcal{TP}}(\mathcal{H})$ is equivalent to a unique element of the form
@@ 5. The operator algebra
We now define a star-algebra structure on $\mathcal{TP}(\mathcal{H}_{\mathbb{R}})$, and eventually on $\overline{\mathcal{TP}}(\mathcal{H}_{\mathbb{R}})$. To distinguish the elements of the algebra from the corresponding elements of the inner product space, we will denote the algebra element $\mathnormal{W}\left(\alpha\otimes_{s}F\right)$ to emphasize its interpretation as a Wick product. Note that this identification differs from the one in Section 3.2. In Remark 5.18 we will interpret the Fock space vector $\alpha\otimes_{s}F$ as the evaluation of the operator $\mathnormal{W}\left(\alpha\otimes_{s}F\right)$ on the vacuum vector in the appropriate GNS representation.
@@
More generally, for $\pi\in\operatorname{\mathcal{P}}_{1,2}(n)$,
@@
The (tensor) contraction is not a bounded operator and so does not extend to the Hilbert space tensor product $\mathcal{H}_{\mathbb{R}}^{\otimes n}$. But if $\pi\in\operatorname{\mathcal{P}}_{1,2}(n,k)$ with $\mathrm{Pair}\left(\pi\right)=\ell$, it is easy to check that
$C_{\pi}:\mathcal{H}_{\mathbb{R}}^{\otimes n}\times\mathcal{H}_{\mathbb{R}}^{\otimes k}\rightarrow\mathcal{H}_{\mathbb{R}}^{\otimes(n+k-2\ell)}$
is a contraction. Indeed, we have the following.
@@
Let $F\in\mathcal{H}_{\mathbb{R}}^{\otimes n}$ and $G\in\mathcal{H}_{\mathbb{R}}^{\otimes k}$. Then
$\left\|C_{\pi}(F\otimes G)\right\|\leq\left\|F\right\|\ \left\|G\right\|.$
Proof.
We will prove the result for the transposition $\pi=(1,n+1)$; the general result is similar with more indices. Let $\left\{\xi_{i}:i\in\Xi\right\}$ be an orthonormal basis for $\mathcal{H}_{\mathbb{R}}$, so that $\left\{\xi_{\mathbf{u}}=\otimes_{k=1}^{n}\xi_{u}(k):\mathbf{u}\in\Xi^{n}\right\}$ is an orthonormal basis for $\mathcal{H}_{\mathbb{R}}^{\otimes n}$. Let $F=\sum_{i,\mathbf{u}}f_{i,\mathbf{u}}\xi_{i}\otimes\xi_{\mathbf{u}}$ and $G=\sum_{j,\mathbf{v}}g_{j,\mathbf{v}}\xi_{j}\otimes\xi_{\mathbf{v}}$. Then by the Cauchy-Schwarz inequality,
$\begin{split}\left\|C_{\pi}(F\otimes G)\right\|^{2}&=\left\|\sum_{i,\mathbf{u},\mathbf{v}}f_{i,\mathbf{u}}g_{i,\mathbf{v}}\xi_{\mathbf{u}}\otimes\xi_{\mathbf{v}}\right\|^{2}\\ &=\sum_{\mathbf{u},\mathbf{v}}\left(\sum_{i}f_{i,\mathbf{u}}g_{i,\mathbf{v}}\right)^{2}\leq\sum_{\mathbf{u},\mathbf{v}}\sum_{i}f_{i,\mathbf{u}}^{2}\sum_{j}g_{j,\mathbf{v}}^{2}=\left\|F\right\|\ \left\|G\right\|.\end{split}$
Next we note that contraction is compatible with the action of the symmetric group.
Lemma 5.4.
Let $\alpha\in S_{0}(n)$, $\sigma\in S(n)$, $F\in\mathcal{H}_{\mathbb{R}}^{\odot n}$ and $\pi\in\mathcal{P}_{1,2}(n)$, with $\left|\mathrm{Pair}\left(\pi\right)\right|=\ell$. Then
@@
$\tilde{\sigma}=P^{[0,n]\setminus\mathrm{supp}\left(\tilde{\pi}\right)}_{[0,n-2\ell]}\sigma|_{[0,n]\setminus\mathrm{supp}\left(\tilde{\pi}\right)}\in S(n-2\ell).$
@@
Then $T$ is also well-defined on $\mathcal{TP}(\mathcal{H}_{\mathbb{R}})$. Note that we cannot in general extend it to $\overline{\mathcal{TP}}(\mathcal{H}_{\mathbb{R}})$.
@@
For $\alpha\in S_{0}(n)$ and $F\in\mathcal{H}_{\mathbb{R}}^{\odot n}$,
$\begin{split}\mathnormal{W}\left(\alpha\otimes F\right)=\mathnormal{T}\left(e^{-\mathcal{L}}(\alpha\otimes F)\right)&=\sum_{\ell=0}^{\infty}(-1)^{k}\frac{1}{\ell!}T\left(\mathcal{L}^{\ell}(\alpha\otimes F)\right)\\ &=\sum_{\pi\in\operatorname{\mathcal{P}}_{1,2}(n)}(-1)^{\left|\pi\right|}\mathnormal{T}\left(C_{\pi}(\alpha\otimes F)\right).\end{split}$
@@
The multiplication on $\mathcal{TP}(\mathcal{H}_{\mathbb{R}})$ is well defined, and
$\mathnormal{T}\left(\alpha\otimes_{s}F\right)\mathnormal{T}\left(\beta\otimes_{s}G\right)]^{\ast}=\mathnormal{T}\left(\beta\otimes_{s}F\right)^{\ast}\mathnormal{T}\left(\alpha\otimes_{s}G\right)^{\ast}.$
@@
Let $\alpha\in S_{0}(n)$, $\beta\in S_{0}(k)$, $F\in\mathcal{H}_{\mathbb{R}}^{\odot n}$, $G\in\mathcal{H}_{\mathbb{R}}^{\odot k}$, $\sigma\in S(n)$, $\tau\in S(k)$. Embed $\alpha,\beta$ in $S_{0}(n+k)$ and $\sigma,\tau$ in $S(n+k)$ as in Notation 3.2, and denote $\rho=\sigma\sigma_{n,k}^{-1}\tau\sigma_{n,k}\in S(n+k)$. Then
@@
Let $q\in\mathcal{Z}\setminus\left\{0\right\}$.
  • $\mathcal{N}_{vs,q}$ is an ideal for multiplication in Definition 5.8.
  • For $\pi\in\mathcal{P}_{1,2}(n)$ with $\left|\mathrm{Pair}\left(\pi\right)\right|=\ell$, $C_{\pi}$ maps $\mathcal{N}_{vs,q,n}$ to $\mathcal{N}_{vs,q,n-2\ell}$.
Consequently, $\mathcal{L}$ is defined as a map on $\mathcal{TP}_{q}(\mathcal{H}_{\mathbb{R}})$, and $\mathcal{L}_{n,k}^{(\ell)}$ on the appropriate quotient. Also, $\mathnormal{T}\left(\eta\right)$ is well-defined for $\eta\in\mathcal{TP}_{q}(\mathcal{H}_{\mathbb{R}})$.
@@
We obtain an extension of the familiar factorization property of Hermite polynomials. Compare also with Proposition 5.17.
@@
Proof.
Use the preceding proposition recursively, and note that a contraction which pairs mutually orthogonal vectors produces a zero term. ∎
The next proposition is the analog of the property that ordinary Hermite polynomial satisfies the differential equation
$xH_{n}^{\prime}-2H_{n}^{\prime\prime}=nH_{n}.$
Note that the Hermite polynomials in Section 2.4 have non-standard normalization.
@@
    for $\eta\in\mathbb{C}[S_{0}(n)]$ and $F\in\mathcal{H}_{\mathbb{R}}^{\odot n}$. Then for such $\eta$,
@@
In particular, anything of lower degree is in the sum of eigenspaces with eigenvalues $0,1,\ldots,n-1$. It follows that specifying the leading term of an eigenfunction with a given eigenvalue determines it. ∎
Notation 5.14.
If $C_{\pi}(\eta\otimes F)$ is a scalar multiple of (0){(0)} (which is the identity for the algebra), we will identify it with a scalar.
@@
Define a unital linear functional $\varphi$ on $\bigoplus_{n=0}^{\infty}\mathbb{C}[S_{0}(n)]\otimes\mathcal{H}_{\mathbb{R}}^{\odot n}$ by
$\varphi\left[\mathnormal{W}\left({(0)}\right)\right]=1,\quad\varphi\left[\mathnormal{W}\left(\alpha\otimes F\right)\right]=0$
for any $\alpha\in\mathbb{C}[S_{0}(n)]$, $F\in\mathcal{H}_{\mathbb{R}}^{\odot n}$, and $n\geq 1$. Then
@@
    In particular, $\varphi$ is well-defined on the quotient $\mathcal{TP}_{q}(\mathcal{H}_{\mathbb{R}})$.
@@
    For $q\in\mathcal{Z}$, $\left\{A:\varphi\left[A^{\ast}A\right]=0\right\}=\mathrm{Span}\left(\mathnormal{W}\left(\zeta\right):\zeta\in\mathcal{N}_{vs,q}\right)$, and $\varphi$ is faithful on $\mathcal{TP}_{q}(\mathcal{H}_{\mathbb{R}})$.
@@
Moreover, each cycle of
@@
intersects $[0,n]$. A calculation shows that its restriction to $[0,n]$ is $\sigma^{-1}\alpha\sigma\beta^{-1}$. Therefore
$q^{n-\operatorname{\mathrm{cyc}}_{0}(\sigma^{-1}\sigma_{n,n}\sigma(\beta^{-1}\cup\alpha))}=q^{n-\operatorname{\mathrm{cyc}}_{0}(\sigma^{-1}\alpha\sigma\beta^{-1})}=\chi_{q}(\sigma^{-1}\alpha\sigma\beta^{-1})$
@@
Consequently, for $q\in\mathcal{Z}$, the star-algebra structure extends to $\overline{\mathcal{TP}}(\mathcal{H}_{\mathbb{R}})$ and $\overline{\mathcal{TP}}_{q}(\mathcal{H}_{\mathbb{R}})$.
@@
$\left|\operatorname{\mathcal{P}}_{1,2}(n,k)\right|$ is sequence A086885 in OEIS. Choosing the locations of the left and right endpoints of $\ell$ pairs, and $\ell!$ ways to pair them, gives an easy estimate
$\left|\operatorname{\mathcal{P}}_{1,2}(n,k)\right|=\sum_{\ell=0}^{\min(n,k)}\ell!\binom{n}{\ell}\binom{k}{\ell}$.
Thus we get a rough estimate
$\left|\operatorname{\mathcal{P}}_{1,2}(n,k)\right|=\sum_{\ell=0}^{\min(n,k)}\frac{n!}{(n-\ell)!}\binom{k}{\ell}\leq n^{k}2^{k}.\qed$
@@
For $\alpha_{i}\in S_{0}(n_{i})$, and $F_{i}\in\mathcal{H}_{\mathbb{R}}^{\otimes n_{i}}$,
@@
Here $\operatorname{\mathcal{P}}_{2}(n_{1},\ldots,n_{k})$ are the inhomogeneous pair partitions [11]. That is, denote by $I(n_{1},\ldots,n_{k})$ the interval partition of $n_{1}+\ldots+n_{k}$ with blocks of sizes $n_{1},\ldots,n_{k}$. Then $\pi\in\operatorname{\mathcal{P}}_{2}(n_{1},\ldots,n_{k})$ if $\pi\in\operatorname{\mathcal{P}}_{2}(n_{1}+\ldots+n_{k})$ and the elements of each pair in $\pi$ belong to different blocks of $I(n_{1},\ldots,n_{k})$.